Sol Summary – May 9th

Crew 261 Sol Summary Report 09-05-2023

Sol 9

Summary Title: Solidarity as a team to solve problem

Author’s name: Aline Decadi, XO

Mission Status: nominal

Sol Activity Summary: Today took place 2 EVAs:
– EVA-13 with the following 5 crew members: Aline Decadi (XO) (EVA Lead), Erin Kennedy (Crew Roboticist), Julien Villa-Massone (Crew Engineer), Cecile Renaud (GreenHab Officer), Kris Davidson (Journalist) (Not In Sim – Not Wearing Suit).

– EVA-14 with the following 2 crew members: Cecile Renaud (GreenHab Officer) (EVA Lead), Julien Villa-Massone (Crew Engineer).

EVA-13 successfully conducted a photo shoot walking outside Hab with suits on.

EVA-14 conducted the rescue of Julien’s Drone successfully + Fly Adapa Drone to get footage, driving with the 2 Opportunity and Curiosity rovers on Cow Dung Road north to Lith Canyon.

The crew spent a part of the afternoon to inspect the toilet due to a bad smell looking for a potential leak, clean it up and fix it. The crew spent the rest of the afternoon: harvesting tomatoes (orange+red), thyme, mint and sage; talk about the Mongolian project MARS-V (creation a laboratorial environment like Mars) advised by Dr Robert Zubrin; have dinner together and prepare the next day’s experiments and EVAs.

Look Ahead Plan: we will conduct 2 EVAs tomorrow: EVA-15 in the morning for engineering testing of drones including fly ADAPA drone to get footage; EVA-16 in the afternoon for the engineering testing of Atmosphinder, according to mission request approval.

Anomalies in work: None.

Weather: Pleasant and sunny. Very windy in the afternoon.

Crew Physical Status: Nominal.

EVA: EVA-13 outside the Hab; EVA-14 on Lith Canyon.

Reports to be filed: HSO Report, EVA Report, Operations Report, Journalist Report, Pictures of the day, GreenHab Report.

Support Requested: None.

Operations Report – May 9th

Crew 261 Operations Report 09-05-2023 Update #1
SOL: 9
Name of person filing report: Julien Villa-Massone
Non-nominal systems:
· Remote observatory
· 1 Helmet, 1 headset
· Some electric outlets in the kitchen
Notes on non-nominal systems: Failed helmet and headset are set aside and labeled as such.
Electric outlets failed on the sink side of the kitchen. Basic troubleshooting consisting of checking breakers did not solve the problem. A cord extender has been installed to temporarily replace these outlets.
Spirit rover used: No
Hours: (before EVA) 220.8
Beginning charge: (Before EVA) 100
Ending charge: (On return from EVA, before recharging) N/A
Currently charging: Yes
Opportunity rover used: Yes
Hours: 125.6
Beginning charge: 100
Ending charge: 68
Currently charging: Yes
Curiosity rover used: Yes
Hours: 236.4
Beginning charge: 100
Ending charge: 59
Currently charging: Yes
Perseverance rover used: No
Hours: 263.1
Beginning charge: 100
Ending charge: N/A
Currently charging: Yes
General notes on rovers: Drove to Lith Canyon. Nominal operation
Summary of Hab operations: Nominal day
WATER USE: rate of 40 gallons / sol today. Trend since mission start of 33.5 gal/day means we will run out of water 2.2 days prior to mission end. Thanks to water resupply, we will now finish the mission with 70 gallons extra if usage continues as per current trend.
Water (static tank): 268 gallons remaining
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: yes. A leak developed due to tank overflow when Crew Engineer was in long EVA. Mission Support was extremely helpful in addressing with the situation. 3 crew members also helped with the process. Toilet issue is now resolved.
Summary of internet: Nominal
Summary of suits and radios: Nominal
Summary of GreenHab operations: Nominal.
WATER USE: 13 gals total
Heater: Off
Supplemental light: Off
Harvest: 755g of tomatoes (orange+red), thym (10g), mint (12g), sage (10g)
Summary of ScienceDome operations: Nominal.
Dual split: Off
POWER report: Nominal. All MPPTs now functional.
Summary of RAM operations: Nominal
Summary of Tunnels maintenance: No zip-ties added.
Summary of any observatory issues: Nothing to report
Summary of health and safety issues: Nominal.
Questions, concerns and requests to Mission Support: Thank you for your support today.

Journalist Report – May 9th

Crew 261 Journalist Report 09-05-2023

Author: Kris Davidson, Crew Journalist

Recently, Commander Burk aptly stated over dinner, "change is the only constant." This time-worn philosophical adage rings true across every facet of existence, from the ceaseless spin of the cosmos to the relentless progression of our own aging bodies. Nowhere does this resonate more than in the realm of science, where change is not just an observer, but the main event itself.

Mars, formed about 4.6 billion years ago, is a testament to the power of change. The planet transitioned from a warm, wet environment with a thick atmosphere and intense volcanic activity to a cold, dry world when it lost its magnetic field about 4 billion years ago. This led to atmospheric erosion by solar wind, and today, Mars’ surface is dotted with impact craters and adorned with polar ice caps composed of water and carbon dioxide. The evolution from an Earth-like planet to its current inhospitable state has scientists intrigued about the possibility of past life on Mars.

The region in Utah surrounding the Mars Desert Research Station (MDRS), also bears the imprints of change. Once submerged beneath prehistoric seas, it saw layers of sandstone, siltstone, and limestone being deposited. The formation and subsequent rifting of the supercontinent Pangea led to uplift and erosion, creating desert landscapes and dinosaur fossil-laden rock formations. Some of these fossils were discovered by MDRS Crew #1, which included Dr. Zubrin!

Transatlantic Mars Crew 261’s stay at the station is already transformative. The portraits taken on Sol 9 on EVA 13 reveal a newfound ease with the Martian environment, the spacesuits, and each other. We may leave Mars by the week’s end, but the experience leaves an indelible mark on us.

Changes and progress are evident in our experiments and projects as well. The crew roots for greenhab officer Cécile Renaud’s young tomato plant growing in Martian regolith stimulant. We’re learning and observing through crew roboticist Erin Kennedy’s ongoing Atmosphider tests, HSO Audrey Derobertmasure’s collecting of vascular aging data, and Julien Villa-Massone’s hab metrics project. James Burk and Aline Decadi continue the vital work of education through Mars VR and other outreach efforts.

The grand narrative of science affirms: change is the only constant. Nothing in our universe remains static. Science, in its quest for understanding, chronicles the dynamic rather than cataloging the static. It documents change, bears witness to the flux, and celebrates the ongoing transformation that underpins all existence.

GreenHab Report – May 9th

Crew 261 GreenHab Report 09-05-2023
GreenHab Officer: Cécile Renaud
Environmental control: Door open from 6:45 to 19:00
Average temperatures: 60.2°F at 06:35, 89.8°F at 12:20, 90.7°F at 16:20, 86.6 at 17:00
Hours of supplemental light: N/A
Daily water usage for crops: 4 gal at 07:15 + 6 gal at 19:00
Daily water usage for research and/or other purposes: 3 gal (including watering solution for biostimulant exp)
Water in Blue Tank 180 gallons (considering the 120gal resupply)
Time(s) of watering for crops: 07:00 and 19:00
Changes to crops: N/A
Narrative: N/A
Harvest: (include which crop and mass in grams) : 755g of tomatoes (orange+red), thym (10g), mint (12g), sage (10g).
Support/supplies needed: None

EVA Report – May 9th

Crew 261 EVA Report 09-05-2023

EVA # 13

Author: James Burk (Commander)

Purpose of EVA: Photo shoot outside Hab with suits on.

Start time: 6:45 AM

End time: 7:50 AM


Overview of EVA
Our crew suited up and went outside to capture several portraits of the crew at dawn, while the lighting was different from daytime, for the purposes of public outreach photos.

Issues Experienced


Destination: Immediate vicinity of main Hab.

Coordinates (use UTM WGS 84): 518400 E, 4250900 N

James Burk (Commander) (EVA Lead), Julien Villa-Massone (Crew Engineer), Kris Davidson (Journalist) (Not In Sim – Not Wearing Suit)

Road(s) and routes per MDRS Map: Walk outside of Hab.

Mode of travel: Walking

Crew 261 EVA Report 09052023

EVA # 14

Author:  James Burk (Commander)

Purpose of EVA: Drone Rescue of Julien’s Drone + Fly Adapa Drone to get footage.

Start time: 6:45 AM

End time: 7:50 AM


Overview of EVA
The EVA team successfully retrieved the lost drone with no issue.  They then proceeded to Lith Canyon and we lost comms for most of their time in that region.  They were able to stop and view (from outside the perimeter) the special region, and then proceed to Lith Canyon where they spent time flying a drone to capture the majestic scenery.  On the way back, they stopped near Pooh’s Corner to place a Meshtastic device that was requested by the remote EVALink team, so that we can test the idea of “breadcrumbs” which can extend the adhoc mesh network that the individual devices form.  Data was gathered and much drone footage was taken from the Adapa drone, and Julien’s DJI drone.

Issues Experienced
• [Technical] The crew found some inaccuracies on the exact road routes near Lith Canyon; the road indicated on our map is hard to find, and there is another, more prominent road, that is not indicated on our map.

• According to the EVA team, the use of one wheel drive on flat ground seemed to extend the rover’s range.  The benefits of this were discussed by the crew.
• The Special Region on the map is not well marked, and it would be easy for somebody to get lost and to enter the area.  Our crew made sure not to enter the area.
• We recommend use of GPS to navigate near and in Lith Canyon (see above issue) because roads are not easy to find and follow.


1. Location of Julien’s Downed Drone
2. Lith Canyon

Coordinates (use UTM WSG 84):
1. 518908 E, 4251236 N
2. 519000 E, 4256600 N

EVA Participants: 2: Cecile Renaud (GreenHab Officer) (EVA Lead), Julien Villa-Massone (Crew Engineer)

Road(s) and routes per MDRS Map: Take Cow Dung Road north to Lith Canyon

Mode of travel: Rovers (Curiosity & Opportunity)

Copyright © The Mars Society. All rights reserved. | Main Site